Sharon is a manufacturer's representative selling office furniture directly to businesses. She receives a monthly salary of $2000 plus a 2.2% commission on sales exceeding her quota of $150,000 per month.
A. What are her earnings for a month in which she has $227,000 in sales?
B. If her average monthly sales are $235,000, what straight commission rate would generate the same average monthly earnings as her current basis of remuneration?

Answer by checkley71(8403)

2000+.022(227000-150000)=2000+.022(77000)=2000+1694=$3694 monthly salary when she sells $227,000.
2000=235000x
2000/235000=x
x=.0085 or .85% commission whem she sells $235,000 to get her base salary of $2000.